### v2

# Three-dimensional quantum criticality

Abstract
Quantum critical phenomena are widely studied across various materials families, from high temperature superconductors to magnetic insulators. They occur when a thermodynamic phase transition is suppressed to zero temperature as a function of some tuning parameter such as pressure or magnetic field. This generally yields a point of instability – a so-called quantum critical point – at which the phase transition is driven exclusively by quantum mechanical fluctuations. Here we show that the heavy fermion metamagnet UTe2 possesses a quantum phase transition at extreme magnetic field strengths of over 70 T. Surprisingly, rather than terminating at one singular point, we find that the phase boundary is sensitive to magnetic field components in each of the three cartesian axes of magnetic field space. This results in the three-dimensional transition surface being bounded by a continuous ring of quantum critical points, the locus of which forms an extended line of quantum criticality – a novel form of quantum critical phase boundary. Within this quantum critical line sits an intensely field–resilient superconducting state in a striking toroidal shape, persisting to fields over 70 T. We model our data by a phenomenological free energy expansion, and show how a three-dimensional quantum critical phase boundary – rather than a more conventional singular point of instability – anchors the remarkable high magnetic field phase landscape of UTe2.
At absolute zero temperature classical phases of matter are completely frozen in place. By contrast, in emergent electronic phases quantum zero-point motion allows for fluctuations to persist down to temperature $T=$ 0 K. For example, when the critical end point of a first-order transition is suppressed to zero temperature, a quantum phase transition [1] occurs at a quantum critical end point (QCEP) [2]. The enhancement of magnetic or density fluctuations close to such a transition can provide the pairing glue for unconventional superconducting phases to condense [3]. For example, $d$-wave superconductivity often accompanies quantum phase transitions involving antiferromagnetism [4, 5], whereas proximity to ferromagnetic ordering may promote the formation of exotic $p$-wave superconductive phases [6, 7, 8]. In a magnetic field the spin degree of freedom of $p$-wave superconductors can yield complex multi-component superconducting phase diagrams [9, 10]. In some rare cases, such as the ferromagnetic superconductor URhGe, the application of a magnetic field H has been found to induce superconductivity close to a high-$H$ QCEP [11, 12, 13].
The paramagnetic metal UTe2 is another intriguing superconducting material. UTe2 possesses a remarkably rich phase diagram, with up to six distinct superconductive phases having been reported under various conditions of magnetic field tilt angle, temperature and pressure [14, 15, 16, 17, 18, 19]. At $\mu_{0}H$ = 0 T a superconducting ground state condenses at $T_{\text{c}}=$ 2.1 K [20, 21, 15, 22]. Despite the absence of ferromagnetism [23], this superconductive state exhibits multiple characteristics of odd-parity $p$-wave pairing [20, 15, 24] analogous to its ferromagnetic uranium-based cousins UGe2 [6], URhGe [7] and UCoGe [8]. The rich multidimensional phase space of UTe2 is further complicated by the presence of charge and pair density wave orderings that have been reported [25, 26] at low-$H$, while at high-$H$ a first-order metamagnetic (MM) transition to a field-polarised paramagnetic (PPM) state occurs for H $\parallel b$, the hard magnetic direction [27].
High field superconductivity in UTe2
Of this menagerie of exotic emergent phases, the focus of the present study is the highest-$H$ superconducting phase found to date in UTe2, which is located at extremely high field strengths starting at $\mu_{0}H\approx$ 40 T and persisting to $\mu_{0}H\gtrapprox$ 70 T. Prior studies have reported this high-$H$ superconductivity to be entirely confined inside the PPM state, nestling against the PPM phase boundary within a narrow angular range for H tilted in the crystallographic $b-c$ plane [18, 28, 16, 29, 30, 31]. Recent measurements have also confirmed its existence for a component of H tilted away from the $b-c$ plane towards the $a$-axis [32] – but again, this exotic high-$H$ superconductivity was found exclusively inside the PPM state. The MM transition into the PPM state has been reported to be strongly first-order at low temperatures for the orientations measured so far [27, 29, 33, 16, 31, 30, 34].
[…]
In combination, our high-$H$ measurements (i) identify that an extended line of QCEPs bounds the surface of first-order MM transitions in three-dimensional H-space at the remarkably high magnetic field scale of over 70 T, (ii) anchor the field–resilient superconductivity in UTe2 close to this boundary, and (iii) reveal that in a narrow angular range close to the QCL the onset of field–resilient superconductivity extends down to fields that lie far below the MM transition field.
Quantum criticality at intense magnetic fields
The highest-field superconducting state in UTe2 is anchored to a quantum critical line associated with a field-induced metamagnetic transition (Figure 1). The key finding of our study is that rotating the orientation of H away from H $\parallel b$, for all directions, the critical end point of the MM transition is suppressed in temperature, eventually reaching 0 K. This thereby forms an extended line (the QCL, labelled in Fig. 1b) composed of QCEPs, demarcating a quantum phase transition at the termination of the PPM phase. Within the QCL sits the high-$H$ superconducting state (labelled SC3), forming a toroidal shape when plotted on Cartesian field axes (Fig. 1a).
We map this evolution of the UTe2 MM transition through contactless resistivity measurements by the proximity detector oscillator (PDO) technique (see Appendix A for details) in pulsed magnetic fields. In Fig. 2a we plot the PDO frequency shift for rotations in the plane that connects the $a$-axis ($\theta_{a}=$ 90°$\degree$, $\theta_{b-c}=$ 0°$\degree$) to the $b-c$ plane at $\theta_{b-c}=$ 17°$\degree$, $\theta_{a}=$ 0°$\degree$; panel (c) shows similar measurements for rotations between the high symmetry directions $b$ and $ac$, i.e. [010]–[101], and from $b$ to $a$, [010]–[100]. For H aligned close to the $b$-axis (i.e. at low values of $\theta_{a}$, $\theta_{b-a}$ and $\theta_{b-ac}$) at low temperatures the MM transition in the contactless resistivity is sharp and first order-like in character, consistent with prior studies [18, 36, 38, 39, 30, 40].
Tilting H away from $b$ increases the magnetic field strength $H_{m}$ at which the transition occurs. As $H_{m}$ increases, the sharpness of the transition smoothly decreases. This softening of the transition with rotation angle at low temperatures is very similar to the behavior for H $\parallel b$ at elevated temperatures plotted in Fig. 2d. At $T=$ 1.5 K the MM jump in PDO frequency is very sharp, exhibiting a large discontinuity in the derivative of the signal with respect to field. The critical end point of UTe2’s MM transition for H $\parallel b$ is reached at $T\approx$ 10 K; above this temperature the transition evolves into a broad crossover [16]. In our PDO measurements at $T=$ 9.0 K for H $\parallel b$ there is still a sharp jump at the transition – by contrast, close to the QCL at $\theta_{b-ac}=$ 42°$\degree$ the transition is markedly softer at $T=$ 1.5 K than it was at 9.0 K for H $\parallel b$.
This suppression of the size of the MM transition is even more apparent in our highest field measurements. In Fig. 2 we plot rotations from $b$ towards $a$ and from $\theta_{b-c}=$ 17°$\degree$ towards $a$, while in Fig. 3 we show the evolution for $\theta_{b-c}=$ 7°$\degree$ towards $a$ and from $b$ towards $c$. In Fig. 3c we show how these measurements enabled us to map the shape of the QCL in three-dimensional H-space. For each rotational plane probed by our experiments, we identify the location of that plane’s QCEP as the angle at which the size of the discontinuity in the derivative of the PDO signal tends to zero, thereby indicating the suppression of the first-order MM phase boundary towards zero temperature. These locations are marked with green stars in Fig. 1 and Fig. 3c. The locus of these QCEPs forms the QCL in three-dimensional H-space.
Condensation of SC3 before the metamagnetic transition
In the ferromagnetic superconductor URhGe, H–induced superconductivity is found in close proximity to a QCEP at $\mu_{0}H\approx$ 10 T, identified by the suppression of a first-order MM transition to zero temperature [12]. Our finding of a similar phenomenon in UTe2 – but in this case exhibiting the remarkable geometric property of quantum criticality stretched along an extended line in H-space at the intense field strengths of 50 T $<\mu_{0}H\lessapprox$ 80 T – motivates a closer examination of SC3 close to the QCL, to which we now turn.
Figure 4 shows contacted resistivity measurements by the standard four-terminal technique for a high quality UTe2 specimen (with a residual resistivity ratio of 420 – see Fig. 9). Up to the maximum applied field of the magnet system utilized for this suite of measurements, of 41.5 T, for incremental magnetic field tilt angles rotating from $b$ towards $ac$ at $T=$ 0.4 K we suddenly access the SC3 state at $\theta_{b-ac}=$ 17.5°$\degree$. In Fig. 4b we define the purple SC3 region by the observation of zero resistivity $\rho$ (in the low current limit). Interestingly, we observe a broadening of the onset of the SC3 phase for successively higher rotation angles $\theta_{b-ac}$, identified as the turning over of $\rho(H)$ from a positive to a negative gradient in $H$ and tracked by the dashed black line in Fig. 4a. This onset extends down to a minimum magnetic field of $\approx$ 30 T at $\theta_{b-ac}\approx$ 34°$\degree$, before progressing back up to higher fields for further inclination in the $b-ac$ plane.
This wide magnetic field domain of negatively sloped magnetoresistance exhibits a highly non-linear relationship between excitation current $I$ and voltage $V$. Whereas standard Ohmic behavior is observed for all measured temperatures at 25 T when the sample is in the normal state, at 41 T a strong deviation is observed at low temperatures (Fig. 4e,f), indicating the likely presence of exotic vortex dynamics [41]. In Figure 5 we plot the exponent $\alpha$ fitted to $V\propto I^{\alpha}$ to illustrate this departure from Ohmic behavior at low temperatures for $\mu_{0}H\gtrapprox$ 30 T. This non-linearity between $I$ and $V$ clearly extends down far below $H_{m}$. We therefore identify a broad range of the UTe2 high field phase diagram, for H tilted close to the QCL, to be occupied by the onset of the SC3 phase.
For H $\parallel b$, the first-order MM transition to the PPM is observed at $\mu_{0}H_{m}\approx$ 34.5 T in the zero temperature limit [27, 29, 42, 36]. As discussed above, on warming the critical end point is crossed at $T\approx$ 10 K and the transition then evolves into a broad crossover, characterized by e.g. a maximum in $\rho(H)$, as the PPM gradually melts and $H_{m}$ decreases until $T\approx$ 30 K where $\mu_{0}H_{m}$ reaches 0 T [15, 16, 27, 29, 34]. Accordingly, we identify the location of $H_{m}$ for each temperature as the field strength at which $\nicefrac{{\partial\rho}}{{\partial H}}=0$, as shown in Figure 6. We schematically plot the corresponding phase landscapes for H $\parallel b$ (i.e. $\theta_{b-ac}=0\degree$) and $\theta_{b-ac}=35\degree$ in Fig. 4c. Strikingly, upon warming above $T_{c}^{SC3}$ at $\theta_{b-ac}=35\degree$, we do not observe a signature of the PPM state in our accessible field range of $\mu_{0}H\leq$ 41.5 T until $T\approx 15$ K (Fig. 6).
Furthermore, we performed a series of rotations in the $b-ac$ plane at 6 K – to suppress all superconducting states and therefore solely probe the PPM phase – to map out the boundary of the MM transition in this plane, which we plot as red circles in Fig. 4b (see Fig. 10 for raw data). The MM transition clearly bisects a substantive region of SC3. Therefore, our measurements have uncovered that a portion of the SC3 phase – and its broad onset region – extends considerably outside the PPM state in the $b-ac$ plane as H is tilted close to the QCL, as we depict in Figure 7. This is in stark contrast to prior measurements of SC3 in the $b-c$ plane, comparatively far from the QCL, that observed SC3 exclusively within the PPM host-state [18, 28, 16, 29, 30, 31, 40].
Landau theory of quantum criticality in three dimensions
The SC3 phase of UTe2, which persists to $\mu_{0}H\gtrapprox$ 70 T despite its relatively modest $T_{c}\sim$ 2 K, has posed a formidable theoretical challenge to our understanding of high-$H$ superconductivity [15, 16]. As prior high-$H$ measurements in the $b-c$ plane only observed SC3 inside the PPM, the field-induced compensation of magnetic exchange [43] was proposed as a possible microscopic origin [31]. However, our finding of the zero resistance SC3 state extending outside the PPM (Fig. 4) – accompanied by an extended field region of the onset of SC3 even further away from the MM phase transition line – argues strongly against this scenario being the sole mechanism.
To shed light on the origin of the QCL, we model the MM transition using a Landau model, including terms up to sixth order in the magnetic order parameter $\mathbf{M}$, the magnetization:
$\mathcal{F}[\mathbf{M}](\mathbf{H})$ $=\frac{1}{2}\chi^{-1}_{i}M_{i}^{2}+\frac{1}{4}\beta_{ij}M_{i}^{2}M_{j}^{2}$ (1)
$+\frac{1}{6}\delta_{ijk}M_{i}^{2}M_{j}^{2}M_{k}^{2}-\mathbf{M\cdot H}$ (2)
where $i,j,k=a,b,c$ and $\chi^{-1}_{i},\beta_{ij}$, and $\delta_{ijk}$ are model parameters (see Appendix B[44, 45, 38]. The metamagnetic phase transition occurs when the free energy has two local minima at $\mathbf{M}=\textbf{M}_{0}$ and $\mathbf{M}=\mathbf{M}_{*}$ corresponding to the normal paramagnetic and PPM phases, respectively. With $\beta_{bb}<0$, the $H_{b}$ component of the external field tends to lower the minimum at $\mathbf{M}_{*}$ faster than the minimum at $\mathbf{M}_{0}$, resulting in the MM transition when $\mathcal{F}(\mathbf{M}_{*})=\mathcal{F}(\mathbf{M}_{0})$, which generically defines a 2D surface in the 3D space of $\mathbf{H}$.
In general, the model parameters are temperature dependent, such that as temperature increases the two minima merge at a critical end point. For simplicity, here we neglect this temperature dependence but consider the effects of $H_{a}$ and $H_{c}$. Due to the mixed terms $\beta_{ij}$ and $\delta_{ijk}$, both $H_{a}$ and $H_{c}$ (which induce $M_{a}$ and $M_{c}$) in effect indirectly modify the $\chi_{b}$ and $\beta_{bb}$ coefficients and so act in a similar way to the temperature. As a result, large $H_{a}$ and $H_{c}$ tend to move $\mathbf{M}_{0}$ and $\mathbf{M}_{*}$ towards each other, until they merge at the critical end line (at finite temperature) or the QCL (at zero temperature) bounding the MM phase transition surface.
We confirm this general picture by numerically minimizing the free energy of Eqn. (2), determining the MM transition surface and the QCL as a function of H, as shown in Figure 8. With the given parameters (see Appendix B) the model matches the experimental data from Fig. 1 reasonably well and captures the main qualitative features. Assuming SC3 is driven by the fluctuations in the vicinity of the QCL, the shape of the QCL explains most of the field direction dependence of SC3, with orbital pair breaking and spin effects potentially accounting for the rest. The QCL fluctuation scenario is further supported by the fact that, analogous to the case of ferromagnetic uranium-based superconductors such as URhGe [12, 13, 45], SC3 should be expected to occur on both sides of the QCL as we observe experimentally (Fig. 4).
Our finding of a three-dimensional quantum critical phase boundary anchoring the extremely high magnetic field superconducting phase landscape of UTe2 poses the question of what the microscopic properties of the intriguing SC3 state may be. Unlike the lower field SC2 phase – the angular domain of which has been found to exhibit a pronounced sensitivity to crystalline disorder [36] – the SC3 state has been observed even in highly disordered crystals that are too dirty to manifest SC1 [40]. Furthermore, in comparing with prior reports [31, 40] we find that the angular extent of SC3 in the $bc$ rotation plane appears to extend to very similar inclination angles irrespective of crystalline quality (see Figure 17). We speculate that this minght be associated with the preferred orientation of triplet spins in the space of $\textbf{H}_{a,b,c}$, and thus may be indicative of the order parameter symmetry. Complementary thermodynamic measurements – although challenging under such extreme magnetic fields – would be instructive for shedding further light on the remarkable quantum critical phenomena at play in UTe2.
[…]
Figure 2: Suppression of the first-order metamagnetic transition towards zero temperature. (a) Contactless resistivity versus field, measured by the PDO technique (see Appendix A). The sample was mounted at a fixed angle of 17°$\degree$ from $b$ towards $c$, and subsequently rotated towards $a$. At a tilt of 6°$\degree$ away from the $bc$ plane the metamagnetic (MM) transition is still sharp and first-order-like; however, by $\theta_{a}$ = 18°$\degree$ it has markedly softened, and at 20°$\degree$ the quantum critical line (QCL) has been passed. (b) Simulation of magnetization for rotations in the same plane as panel (a), calculated from Eqn. 2 showing how the sharp jump in magnetization is suppressed upon passing the QCL. (c) PDO data for the $b-a$ ([010]–[100], top) and $b-ac$ ([010]–[101], bottom) rotation planes. All curves to $<$ 60 T were measured in a single coil magnet in Wuhan at $T=$ 4.2 K, while those to $>$ 60 T were measured in a dual-coil magnet in Dresden at 1.7 K; note that the Dresden data have been rescaled for ease of comparison, while the nature of the dual-coil magnet system introduces a small artifact at 39 T (see Fig. 15 for details). Like in panel (a), the MM transition is suppressed towards zero temperature as the QCL is approached. (d) Temperature evolution of the MM transition for H $\parallel b$ (top) and at $\theta_{b-ac}=$ 42°$\degree$ (bottom); derivatives are plotted in panel (e). The 9.0 K transition for H $\parallel b$ is considerably sharper than the 1.5 K curve at $\theta_{b-ac}$ = 42°$\degree$, which lies close to the QCL.
Figure 3: Mapping the quantum critical line in three dimensions. (a) PDO data measured in Dresden rotating from $\theta_{b-c}$ = 7°$\degree$ towards $a$ at 1.7 K. Within a one degree inclination from $\theta_{a}$ = 17°$\degree$ in this plane, the QCL is abruptly crossed. (b) PDO data of the MM transition in the $b-c$ plane, measured up to 57 T in Wuhan and up to 80 T in Dresden. The Dresden data has been rescaled by a factor of 1.1 to be comparable with that from Wuhan (see Fig. 18). (c) Positive quadrant of $H_{a}-H_{c}$ showing how the QCEPs were identified from our rotational measurements in high magnetic fields. The size of orange points corresponds to the size of the MM transition.
Figure 4: Spillover of SC3 outside the polarised paramagnetic phase. (a) Electrical resistivity $\rho(H)$ at successive magnetic field tilt angles in the $b-ac$ rotation plane at $T=$ 0.4 K. At $\theta_{b-ac}=$ 17.5°$\degree$ zero resistance is observed over the field interval of 0 T $\leq\mu_{0}H\leq$ 41 T as each of SC1, SC2 and SC3 are successively accessed. (b) The phase landscape of UTe2 plotted radially in $\theta_{b-ac}$. An extended region of the onset of SC3 is observed outside the PPM state, identified by the turning over from positively to negatively sloped $\rho(H)$ marked by the black dashed line in panel (a). The red circles defining the PPM boundary were measured above $T_{c}^{SC3}$. (c) The evolution of the SC1, SC2, SC3 and PPM phases for $\theta_{b-ac}=$ 0°$\degree$ (left) and $\theta_{b-ac}=$ 35°$\degree$ (right). At $\theta_{b-ac}=$ 35°$\degree$ the melting of PPM is not observed at 5 K (marked by a cross – see Fig. 6). This shows that SC3 extends outside the PPM state for H tilted close to the QCL. (d) A comparison of the temperature sensitivity of SC3 at $\theta_{b-ac}=$ 18°$\degree$ (left) and $\theta_{b-ac}=$ 35°$\degree$ (right). On the edge of the SC3 dome at $\theta_{b-ac}=$ 18°$\degree$ the $T_{\text{c}}$ is only $\approx$ 0.5 K, whereas at $\theta_{b-ac}=$ 35°$\degree$ SC3 is much more robust to elevated temperatures. (e) Effective resistivity at $\theta_{b-ac}=$ 35°$\degree$ for incremental excitation currents as indicated. (f) Log-log plot of voltage $V$ versus current $I$ from 0.4-1.8 K at $\theta_{b-ac}=$ 35°$\degree$. Ohmic $I-V$ characteristics are observed in the normal state at $\mu_{0}H=$ 25 T for all currents, whereas at $\mu_{0}H=$ 41 T the $I-V$ behavior is highly non-Ohmic. We therefore identify the turning over of the $\rho(H)$ traces in (a) as marking the onset of SC3, characterized by this strongly non-linear relationship between $I$ and $V$.
Figure 5: Anomalous $I-V$ characteristics of SC3. The exponent $\alpha$ extracted from fitting $V\propto I^{\alpha}$ to isothermal field sweeps at incremental currents, plotted as a function of field at incremental temperatures (left) and as a function of temperature for incremental fields (right). The raw data from which this figure is constructed are plotted in Figure 11.
Figure 6: (a) $\rho(H)$ for H $\parallel b$ at incremental temperatures, with corresponding $\nicefrac{{\partial\rho}}{{\partial H}}$ in (b). The inset shows the derivative of the 28 K curve. (c) $\rho(H)$ at $\theta_{b-ac}$ = 35°$\degree$, with corresponding $\nicefrac{{\partial\rho}}{{\partial H}}$ in (d). These data were used to construct Fig. 4c. In all panels, arrows mark zero $\nicefrac{{\partial\rho}}{{\partial H}}$, identifying the location of the metamagnetic transition. Importantly, zero resistance is observed at $\mu_{0}H\approx$ 40 T at low temperatures for $\theta_{b-ac}=$ 35°$\degree$ even though $\nicefrac{{\partial\rho}}{{\partial H}}$ rises monotonically at $T$ = 5 K, indicating that in the low temperature limit for this orientation the metamagnetic transition field is located significantly above the maximum accessed field strength of 41.5 T for these measurements – even though zero resistance is observed at low temperatures. This shows that the SC3 phase has a region of spillover outside the PPM state.
Figure 7: Construction of the high magnetic field phase diagram. Here we plot the positive magnetic field octant of the schematic phase diagram displayed in Fig. 1a, including data points used to bound the SC1, SC2, SC3 and PPM phases. Each data point corresponds to either a DC field sweep or a pulsed field measurement. The open square point marking the top of SC3 is taken from ref. [31]; all other points are from this study. The SC3 spillover and onset regions from Fig. 4 are also identified.
Figure 8: Numerical simulation of the MM transition and QCL using a Landau free energy expansion. As in Fig. 1, the orange surface corresponds to the surface in $\mathbf{H}$–space at which the first-order MM transition occurs, obtained by interpolation from numerically determined locations of the jumps in the magnetization found by minimizing the free energy of Eqn. (2). The QCL (green points) was obtained by determining the location where the jumps vanish within numerical accuracy. The parameters were chosen as follows: $\chi_{a}^{-1}=36.13$ (easy axis), $\chi_{b}^{-1}=\beta_{cc}=180.7$ (hard axis), $\chi_{c}^{-1}=\beta_{aa}=72.27$, $\beta_{bb}=-688.4$, $\beta_{ab}=0.18$, $\beta_{bc}=-361.3$, $\delta_{bbb}=903.4$, $\delta_{ccc}=1445$, $\delta_{abb}=903.4$, $\delta_{bbc}=1084$, and all others zero.
Acknowledgements.
We are grateful to A. Levchenko, N.R. Cooper, T. Helm, T. Hazra, P. Coleman and J. Schmalian for stimulating discussions. We thank S.T. Hannahs for technical advice and assistance. This project was supported by the EPSRC of the UK (grant no. EP/X011992/1). A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement Nos. DMR-1644779 & DMR-2128556 and the State of Florida. We acknowledge support of the HLD at HZDR, a member of the European Magnetic Field Laboratory (EMFL). The EMFL also supported dual-access to facilities at MGML, Charles University, Prague, under the European Union’s Horizon 2020 research and innovation programme through the ISABEL project (No. 871106). Crystal growth and characterization were performed in MGML (mgml.eu), which is supported within the program of Czech Research Infrastructures (project no. LM2023065). We acknowledge financial support by the Czech Science Foundation (GACR), project No. 22-22322S. T.I.W. and A.J.H. acknowledge support from EPSRC studentships EP/R513180/1 & EP/M506485/1. T.I.W. and A.G.E. acknowledge support from ICAM through US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of the Office of International Science and Engineering and from QuantEmX grants from ICAM and the Gordon and Betty Moore Foundation through Grants GBMF5305 & GBMF9616. D.V.C. acknowledges financial support from the National High Magnetic Field Laboratory through a Dirac Fellowship, which is funded by the National Science Foundation (Grant No. DMR-1644779) and the State of Florida. The work at UW-Madison (D.S.) was financially supported by the National Science Foundation, Quantum Leap Challenge Institute for Hybrid Quantum Architectures and Networks Grant No. OMA-2016136. A.G.E. acknowledges support from the Henry Royce Institute for Advanced Materials through the Equipment Access Scheme enabling access to the Advanced Materials Characterisation Suite at Cambridge, grant numbers EP/P024947/1, EP/M000524/1 & EP/R00661X/1; and from Sidney Sussex College (University of Cambridge).
Appendix A EXPERIMENTAL TECHNIQUES
Sample preparation
High quality single crystals of UTe2 were grown by the molten salt flux method, adapted from ref. [22]. Samples were grown by the methodology given in ref. [35]. Specimens were oriented by single crystal Laue diffractometry, and screened by a combination of residual resistivity, specific heat, and magnetization measurements.
Electrical conductivity measurements
Contacted transport measurements were performed by the four-wire technique using low frequency AC excitations ($<$ 50 Hz) with voltages probed by phase sensitive detection. Electrical contacts were made by spot-welding gold wires onto samples, which were then mechanically secured with low-temperature epoxy.
Contactless electrical conductivity measurements were made by the proximity detector oscillator (PDO) technique [47]. Experiments were performed at the Wuhan National High Magnetic Field Center in pulsed magnetic fields up to 56.7 T, and at the Hochfeld-Magnetlabor Dresden (HLD-EMFL) in a dual-coil magnet to a maximum applied field of 79.7 T. PDO measurements were performed at frequencies of tens of MHz, giving good resolution despite the short duration of a magnetic field pulse, of order 0.1 s (see Figure 15). By tracking the change in frequency $f$ of the PDO circuit [47, 48] as a function of $H$, the change in a sample’s electrical resistivity $\rho$ is captured as
$\frac{\Delta f}{f}\approx-\eta\frac{\delta}{d}\left(\mu_{r}\frac{\Delta\rho}{\rho}+\Delta\chi_{s}\right),$ (3)
for a sample of thickness $d$, where $\mu_{r}=\chi_{s}+1$ and the skin depth $\delta=\sqrt{\frac{2\rho}{\mu_{r}\mu_{0}\omega}}$, where $\omega$ is the excitation frequency [47, 48]. Our PDO circuitry typically rings in the range 28-32 MHz – in Dresden this signal was then mixed down once to $\approx$ 11 MHz and acquired using a Teledyne LeCroye HDO6000B oscilloscope acquiring at 500 megasamples per second. In Wuhan we mixed down twice and acquired at $\approx$ 2 MHz utilizing a JYTEK PXIe-69834 data acquisition card at a rate of 30 megasamples per second.
Appendix B LANDAU THEORY OF METAMAGNETIC TRANSITION AND QUANTUM CRITICAL LINE
The theoretical simulation of magnetization curves in Fig. 2b was obtained by numerically searching for local minima of the free energy in Eq. 2, picking the lowest minimum found. The critical field $\mathbf{H}_{m}(\theta_{b-c},\theta_{a-c})$ at which the MM transition occurs for a fixed field direction was identified by iteratively searching for the largest rate of change in the magnetization. The QCEP magnetic field values $\mathbf{H}_{QCEP}(\theta_{a-c})$ were then obtained by scanning over different field directions starting from the $b$ axis and incrementing the $\theta_{b-c}$ and $\theta_{a-c}$ angles in steps of $1.8\degree$, terminating when the numerically computed slope of the magnetization curve along the scanning direction at the MM transition fell below a threshold (determined by the magnetization slopes above and below the transition). The locus of the points $\mathbf{H}_{m}(\theta_{b-c},\theta_{a-c})$ and $\mathbf{H}_{QCEP}(\theta_{a-c})$ thus obtained were used, respectively, to plot the MM phase transition surface (orange, obtained by interpolation) and the QCL line (green) in Figure 8.
Now we turn to provide more qualitative details of the Landau theory of the metamagnetic (MM) phase transition described by the free energy in Eq. 2 in the main text:
$\mathcal{F}[\mathbf{M}](\mathbf{H})$ $=\frac{1}{2}\chi^{-1}_{i}M_{i}^{2}+\frac{1}{4}\beta_{ij}M_{i}^{2}M_{j}^{2}$ (4)
$+\frac{1}{6}\delta_{ijk}M_{i}^{2}M_{j}^{2}M_{k}^{2}-\mathbf{M\cdot H}.$ (5)
To understand the origin of the critical end point (CEP) and the quantum critical end point (QCEP), we first review a simpler case of a 1D model:
$\mathcal{F}[M](H)$ $=\frac{1}{2}\chi^{-1}M^{2}+\frac{1}{4}\beta M^{4}$ (6)
$+\frac{1}{6}\delta M^{6}-MH.$ (7)
The ground state is found by minimizing the free energy with respect to $M$, which gives
$H=\chi^{-1}M+\beta M^{3}+\delta M^{5}.$
At the MM transition, $\partial_{M}H=0$, which can be solved analytically for $M$:
$M^{2}_{\pm}=\frac{-3\beta\pm\sqrt{9\beta^{2}-20\chi^{-1}\delta}}{10\delta}$
with $M_{\pm}$ being the values of $M$ immediately above and below the MM transition, respectively. In order for the MM transition to exist, we must therefore have $\beta<0$ and $09\beta^{2}-20\chi^{-1}\delta>$ 0, and the MM jump vanishes precisely when $09\beta^{2}-20\chi^{-1}\delta=$ 0. The latter condition thus determines the CEP when the Landau parameters $\chi^{-1},\beta,$ and $\delta$ are functions of some external parameter like temperature $T$:
$09\beta^{2}(T_{CEP})-20\chi^{-1}(T_{CEP})\delta(T_{CEP})=$ 0 .
In the case of UTe2, $T_{CEP}\approx$ 10 K for H $\parallel b$ [27, 34].
Generically, the CEP occurs due to $\beta(T)$ increasing with increasing temperature and eventually changing sign at some temperature $T_{0}>T_{CEP}$. As an example, in itinerant systems, to leading order close to temperature $T_{0}$ at which $\beta(T)=0$, the temperature dependence of the coefficients can be modelled phenomenologically as [44]
$\chi^{-1}_{i}(T)$ $\propto 1+bT^{2}-bcT^{4},$
$\beta_{ij}(T)$ $\propto 1-cT^{2}.$ (8)
When the parameters $b$ and $c$ are functions of another external parameter $\theta$, i.e. the field direction, the corresponding $T_{CEP}$ also becomes a function of $\theta$ and may even vanish for some $\theta_{QCP}$, resulting in a quantum critical point.
To obtain a qualitative understanding of the 3D free energy in Eq. (2), we can rewrite $\mathbf{M}$ in spherical coordinates, which effectively gives
$\mathcal{F}[M,\theta,\phi](\mathbf{H})$ $=\frac{1}{2}\chi^{-1}(\theta,\phi)M^{2}+\frac{1}{4}\beta(\theta,\phi)M^{4}$ (9)
$+\frac{1}{6}\delta(\theta,\phi)M^{6}-MH(\sin\phi\sin\phi_{H}$
$\cdot\cos(\theta-\theta_{H})+\cos\phi\cos\phi_{H})$ (10)
where $\theta$ and $\phi$ are azimuthal and polar angles of $\mathbf{M}$ and $\theta_{H}$ and $\phi_{H}$ are azimuthal and polar angles of $\mathbf{H}$. For fixed $\mathbf{H}$, minimizing over $\theta$ and $\phi$ determines the direction of $\mathbf{M}$, which implicitly determines $\theta(\mathbf{H})$ and $\phi(\mathbf{H})$. Consequently, we can consider $\chi^{-1},\beta$, and $\delta$ to be functions of $\mathbf{H}$, as well as temperature. Strictly speaking, we should add a coefficient in front of the $MH$ term in Eq. 10 that is also a function of $\mathbf{H}$, adding a nonlinearity in $H$ to the problem; however, since the CEP condition is independent of $H$ we can neglect this nonlinearity as long as it is not too strong. Setting the temperature to zero, we then have a QCP condition given by
$09\beta^{2}(\mathbf{H}_{*})-20\chi^{-1}(\mathbf{H}_{*})\delta(\mathbf{H}_{*})=$ 0 .
This equation defines a 2D surface in the 3D space of $\mathbf{H}$, which generically intersects the 2D MM transition surface (determined by $\partial_{\mathbf{M}}\mathbf{H}=0$) in a line, thereby forming the quantum critical line (QCL). Qualitatively, given that $\beta(H_{a},H_{b},H_{c})$ is minimal for $H_{a}=H_{c}=0$, the field components $H_{a}$ and $H_{c}$ act effectively in the same way as temperature as they increase $\beta$ towards zero.
Appendix C SUPPORTING FIGURES
Figure 9: Zero field contacted resistivity versus temperature measured by the four-terminal technique on the same sample as that investigated in DC magnetic fields up to 41.5 T for which data are presented in Figure 4. Current density $j$ was sourced parallel to the $a$-axis. The $T_{\text{c}}$ value of 2.09 K is determined from the resistivity falling to zero (within the resolution of the measurement). Extrapolating the normal state resistivity below 5 K with a quadratic fitting towards zero temperature (red curve) yields a residual resistivity ratio (RRR) of 420, indicative of high crystalline quality [22, 36, 37].
Figure 10: Identification of the spillover of SC3 outside the PPM. (a) DC magnetic field sweeps at $T$ = 0.4 K in the $b-ac$ rotation plane. A subset of these data are presented in Fig. 4a. These data are used to construct the phase diagram in Fig. 4b. (b) DC field sweeps at the elevated temperature of $T$= 6 K – above all critical superconducting temperatures – in order to map the extent of the PPM phase in this rotation plane. The MM transition is observed as a peak in $\rho(H)$ at incremental rotation angles up to $\theta_{b-ac}$ = 27.5°$\degree$; however, at the next measured angle of $\theta_{b-ac}$ = 31.0°$\degree$ no peak is observed up to the highest applied field of 41.5 T. In panel (a) the 0.4 K data at $\theta_{b-ac}$ = 31.0°$\degree$ is plotted in the same blue as panel (b), clearly showing that the zero resistance state of SC3 is accessed even though the MM transition for this field orientation lies at higher field strengths.
Figure 11: Non-Ohmic current–voltage behavior near the QCL. Effective resistivity $\rho$ for incremental currents at isothermal temperature points as indicated. We note that some Joule heating likely occurs for the highest current measurements at 0.5 K (and to a lesser extent for 0.7 K) as seen by the change in slope of $\rho(H)$ in the normal state over the interval 25 T $\lessapprox\mu_{0}H\lessapprox$ 30 T for e.g. the 20 mA curve compared to that at 150 $\upmu$A. However, there is clearly an intrinsic dependence on current for the onset of SC3 between the 150 $\upmu$A and 400 $\upmu$A curves, which exhibit no indication of extrinsic heating effects. Furthermore, at higher temperatures the linear relationship between $V$ and $I$ in the normal state (i.e. at $\mu_{0}H=$ 25 T) holds extremely well for all currents, while even up to 1.8 K at $\mu_{0}H\approx$ 40 T the SC3 onset regime exhibits clearly non-linear $I-V$ behavior. The 1.1 K data are those displayed in Fig. 4e.
Figure 12: Tracking the softening of the metamagnetic transition as the QCL is approached. Contactless resistivity measurements by the PDO technique in the (a) $b-c$, (b) $b-a$ and (c) $b-ac$ rotation planes, with corresponding derivatives in (d-f). All data were collected on the same sample at $T=$ 4.2 K, measured in the Wuhan National High Magnetic Field Center.
Figure 13: Temperature evolution of the metamagnetic transition. PDO data at incremental temperatures for six different magnetic field orientations, as labelled. All data were collected on the same sample, measured in the Wuhan National High Magnetic Field Center.
Figure 14: PDO measurements tracking the angular extent of the SC3 phase up to $\mu_{0}H$ = 57 T at $T$ = 0.7 K in (a) the $b-c$ plane; (b) the $b-ac$ plane; (c) from $\theta_{b-c}$ = 33°$\degree$ rotating towards $a$.
Figure 15: (a) Raw PDO data from a pulsed field sweep using the Dresden dual-coil magnet system. (b) Raw PDO data from a pulsed field sweep using single-coil magnet system in Wuhan. (c) Magnetic field versus time for representative magnetic field pulses performed in Dresden (indigo) and Dresden (ochre). An arrow indicates the field profile of the second coil in Dresden. (d) The same data as in (a) plotted as a function of time. The small artifacts in our Dresden measurements at 39 T stem from the activation of the second coil, and are not intrinsic features of UTe2, unlike the metamagnetic transition.
Figure 16: High field UTe2 phase diagram for magnetic field orientations in the $bc$ and $ba$ rotational planes. The phase boundaries of the SC1, SC2, SC3 and PPM phases as determined by our prior measurements reported in ref. [36] and by those from this study.
Figure 17: Comparison of the angular extent of SC3 in the $b-c$ rotation plane from measurements of electrical conductivity presented in this study along with those reported in refs. [31, 40]. All data are at $T\approx$ 0.5-0.7 K. Orange points represent the transition into the PPM state, while purple and blue symbols indicate the start and end of the SC3 state. Black dashed lines mark the highest angle at which SC3 is still observed; orange dashed lines mark the next subsequent measured angle from each study, where signatures of superconductivity are no longer resolved (only the transition to the PPM phase). Note that two samples are plotted from ref. [40]. The SC3 state seems to terminate in the same range of inclination of 41°$\degree\lessapprox\theta_{b-c}\lessapprox$ 45°$\degree$ in each of these studies.
Figure 18: Extrapolation of the $b-c$ plane metamagnetic transition to its QCEP at high field. Note that Dresden data in Fig. 2 has been rescaled by a factor of 0.57 to compare with Wuhan data, while here we plot data from Dresden measured on a different sample in a different coil that has been rescaled by a factor of 1.1, which gives the same size of MM jump at $\theta_{b-c}$ = 24°$\degree$ (where $\mu_{0}H_{m}=$ 37.4 T). By taking the size of the metamagnetic transition as a function of $\theta_{b-c}$, we fit a linear trendline to extrapolate to the QCEP and plot open green stars at $|\mu_{0}H_{c}|$ = 69 T ($\mu_{0}H_{m}=$ 76.7 T) in Fig. 1.
Figure 19: Comparison of high field PPM and SC3 data measured in this study with those reported in Lewin et al. [32]. Our data are plotted with solid symbols, while those of ref. [32] are given as open symbols. Solid green stars are QCEPs from our measurements, while open green hexagrams are interpreted from the data in ref. [32]. We plot an indicative error bar for the QCEPs interpreted from the data in Lewin et al. [32], as those data were measured below $T_{c}^{SC3}$ therefore it is difficult to definitively determine the location of the QCEP as the domain of SC3 can stretch on either side of the QCL, as we show in Fig. 4. We therefore include these indicative error bars, constructed by spanning the angular region halfway between the last measurement point clearly showing first order–like metamagnetism and the suppression towards 0 K. We find that the dataset reported in Lewin et al. [32] agrees very well with our construction of the UTe2 high magnetic field QCL.
Figure 20: Temperature evolution of the metamagnetic transition for H $\parallel b$. Circular points are from Fig. 6 with triangular points from ref. [49]. The metamagnetic transition field $H_{m}$ decreases monotonically with elevated temperatures. Our observation of monotonically rising $\rho(H)$ (and $\nicefrac{{\partial\rho}}{{\partial H}}$) at $\theta_{b-ac}=$ 35°$\degree$ for $T=$ 5 K in Fig. 6 therefore shows that $\mu_{0}H_{m}$ lies higher than 41.5 T. However, remarkably, at low temperatures we observe zero resistance at $\mu_{0}H<$ 41.5 T (Fig. 4), indicating the spilling over of SC3 beyond the PPM phase.